OpenStudy (linda3):

Find the slope and the y-intercept of the line. y=-3/2x+6

OpenStudy (luigi0210):

\[\LARGE y=mx+b\] m=slope and y-intercept is when x=0.. so the b

OpenStudy (luigi0210):

So pretty much slope=\[\LARGE -\frac{3}{2}\] and intercept is 6

OpenStudy (linda3):

so would that mean slope = -3/2 y-intercept= x=-2/3y-4 ???

OpenStudy (linda3):

nvm... thats not right

OpenStudy (luigi0210):

When looking for an intercept , just plug in 0 for the value. So in this case: \[\LARGE y=-\frac{3}{2}(0)+6\] And anything times 0=0 so: \[\LARGE y=0+6\] Which is just \[\LARGE y=6\] In coordinates: \[\LARGE (0, 6)\]

OpenStudy (linda3):

no.... I wished you told me that sooner D: I got it wrong

OpenStudy (linda3):

lol

OpenStudy (luigi0210):

Awe, sorry. But that's why you confirm it first my dear~ But good luck.

OpenStudy (linda3):

hehehe thank you